OpenStudy (anonymous):

What is the range of 3/x2-4

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

\(f(x) = \dfrac{3}{x^2}-4\) ?

OpenStudy (decentnabeel):

\[\mathrm{Range\:of\:}\frac{3}{x^2}-4:\quad \begin{bmatrix}\mathrm{Solution:}\: & \:f\left(x\right)>-4\: \\\:\mathrm{Interval\:Notation:} & \:\left(-4,\:\infty \:\right)\end{bmatrix}\]

OpenStudy (anonymous):

I think it's \[\mathbb{R}-\left\{ -4 \right\}\] When f(x)=-4, we have \[\frac{3}{x^2}-4=-4\] \[\implies \frac{3}{x^2}=0\]\[\implies 3=0\] which is not true I'm not sure though